OpenStudy (anonymous):

plz HELPP Fan+ Medal 108 is 36% of what number? write and solve a proportion to solve the problem

OpenStudy (xxhomeworkhelperxx):

108 is 36% of what number? so "of" means times 36%=0.36 and x is our unknown number we want to find so we can set it out like this. 0.36 times x = 108 x = 108/0.36 x = ........... (you can do this)

OpenStudy (xxhomeworkhelperxx):

xplanation: I know "of" means times and 36%=0.36 and I make x my unknown, then I can set out an equation with what was given. 0.36 times x = 108 then I can solve for x from there to find out the number.

OpenStudy (michele_laino):

from text of your problem, I read that: \[108=\frac{ 36 }{ 100 }*x\] so, your proportion is: \[108:36=x:100\]